OpenStudy (anonymous):

a plane ticket costs b dollars for an adult and d dollars for a child. Find the total cost of 5 adult tickets and 3 child tickets.

OpenStudy (anonymous):

Well, how much money is an adult ticket?

OpenStudy (anonymous):

it doesn't say, that's really all the information the problem gives me. . .

OpenStudy (anonymous):

If this is the case, then just use the variable as your answer.

OpenStudy (anonymous):

b and d would be the answer pretty much then?

OpenStudy (anonymous):

So, if there is 5 AT you want to buy, and each cost b dollars, then how much money will you spend on ATs?

OpenStudy (anonymous):

5b

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

Yes, correct, and how much would you spend on child tickets?

OpenStudy (anonymous):

5d

OpenStudy (anonymous):

How much does each child ticket cost though?

OpenStudy (anonymous):

ill put the problem again. a plane ticket costs b dollars for an adult and d dollars for a child. Find the total cost of 5 adult tickets and 3 child tickets. so out of the examples you put there all i get to understand now is 5b and 3d

OpenStudy (anonymous):

there's no price, just amount of adults/children and the variables

OpenStudy (baldymcgee6):

thats right @ryuu just 5b+3d

OpenStudy (anonymous):

alrighty! :)!

OpenStudy (anonymous):

Yes, if there is no more information, then most likely, C(cost) is going to be 5b + 3d

OpenStudy (anonymous):

thanks guys :)

OpenStudy (anonymous):

No prob

OpenStudy (anonymous):

fan me mcgee lol

OpenStudy (anonymous):

=P

OpenStudy (baldymcgee6):

haha, but you haven't helped me yet.. :)

OpenStudy (anonymous):

pfft, i wanna send you a msg! XD

OpenStudy (anonymous):

If you need any help, I am open for tutoring :)

OpenStudy (anonymous):

baldy all you did was confirm the answer :P

OpenStudy (anonymous):

that'd really help, i need some more help with like 5 more word problems or a bit more.. no more than 10 though, xD

OpenStudy (anonymous):

Ok lol I can help you here free and if you want I can tutor you every week for a certain amount of time(20$/hour) and I could give you some practice problems and help you understand them, etc. :)

OpenStudy (baldymcgee6):

@chandhooguy are you a teacher or something?

OpenStudy (anonymous):

nah i barely use the site 3 times a week the most, i just need help on a few more cause the word problems feel tricky to me and thats what i have problems with, asides that my packet will be finished after that, :P

OpenStudy (baldymcgee6):

ryuu whats the next question

OpenStudy (anonymous):

@baldymcgee6 Yes

OpenStudy (anonymous):

Ok ryuu, just a suggestion. I can still help though :)

OpenStudy (anonymous):

the product of two numbers is 13. One number is k, what is the other number? I came down to the equation of k + ? = 13

OpenStudy (anonymous):

thanks a lot =)

OpenStudy (anonymous):

Remember, product means Multiplication

OpenStudy (baldymcgee6):

remember product is when you multiply not add

OpenStudy (anonymous):

Ohh!! >_< i keep forgetting.

OpenStudy (anonymous):

then.. k * ? = 13 >?

OpenStudy (baldymcgee6):

so its k*x = 13

OpenStudy (anonymous):

K? = 13 Isolate K

OpenStudy (baldymcgee6):

yeah, instead of using question marks, in math we typically use variables

OpenStudy (anonymous):

alright, so.. divide x to both sides?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

k * x/x = 13/x > ?

OpenStudy (baldymcgee6):

right

OpenStudy (baldymcgee6):

but what is x/x ?

OpenStudy (anonymous):

=> k = 13/x <=?

OpenStudy (anonymous):

Well, you want to isolate x, not k, so you want x = ...

OpenStudy (anonymous):

oh.

OpenStudy (anonymous):

So how do you isolate ONLY x?

OpenStudy (anonymous):

so then.. i would have to divide k and not x? lol

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

alrighty got it down now. =)

OpenStudy (anonymous):

Ok I will be back in 1/2 hour, so hang on or ask bald till then :)

OpenStudy (anonymous):

Lol, okay. He seems busy though =P

OpenStudy (anonymous):

The Police and Kenny Chesney had the two top-grossing North American concert tours in 2007, together generating $204.3million in ticket sales. If Kenny Chesney took in $62.1million leess than The Police, how much did each tour generate?

OpenStudy (anonymous):

i used logic on this but i may be way off.. -->(total) 204.3 - 62.1(kenny chesney)= 142.2 / 2 = (Chesney)>71.1 + 62.1 =(The Police) $133.2million. +71.1 = 204.3(Total) The Police: $133.2million Kenny Chesney: $71.1million Total: $204.3million

OpenStudy (anonymous):

No no, Chesney did not take 62.1 million. We don't know how much he took, but we do know that he took 62.1 million LESS THAN the Police. How you would solve this problem would be by setting up 2 variables, x being how much Chesney got and y being the amount the Police got.So, you know that in all, 204.3 million was sold, and Chesney sold 62.1 million less than the Police. The formula would be... x + y = 204,300,000 .... This is just saying that what they both got is equal to 204.3 million x = y - 62,100,000 ... This is just saying that Chesney got exactly how much the Police got, MINUS 62.1 million Right now, this all should make sense. Just plugin equation 2 into equation 1, and see what you get. If you have any doubts, tell me an I will be glad to answer.

OpenStudy (anonymous):

Excuse my last post. If you do it either way, you get the same answer, but my way is just the 'Algebraic' way of doing it. Sorry about that.